OpenStudy (anonymous):

E-Loan, an online lending service, recently offered 60-month auto loans at 3.9% compounded monthly to applicants with good credit ratings. If you have a good credit rating and can afford monthly payments of $587, how much can you borrow from E-Loan? what is the total interest you will pay for this loan? Was wondering if someone can help me with the formula for this problem and maybe walk me through it :)

OpenStudy (anonymous):

Its not a future value? its a Present Value?

OpenStudy (zzr0ck3r):

\(A=P(1+\frac{r}{n})^{nt}\) where n is the amount of times its compounded yearly. So n for us is 12 \(A=P(1+\frac{0.039}{12})^{12t}\) since you can afford 587 we do this for 60 months so \(587*60=P(1+\frac{0.039}{12})^{12*5}\)

OpenStudy (zzr0ck3r):

I think

OpenStudy (zzr0ck3r):

about 29 thousand dollars.

OpenStudy (zzr0ck3r):

http://www.wolframalpha.com/input/?i=587*60%3Dx%281%2B \frac{0.039}{12}%29^{12*5}

OpenStudy (anonymous):

That sounds correct to me, I got 28989.40 Do you know how to see the interest paid?